# Memory-Efficient Acceleration of Block Low-Rank Foundation Models on Resource Constrained GPUs

### v1

@@ 6 Conclusion
Our optimized BLAST and Monarch routines still lag behind low-rank decompositions in terms of speed, a limitation rooted in their blocked structure, which generates more intermediate outputs. This overhead, however, could be mitigated by future techniques such as intermediate activation quantization, provided accuracy is preserved and target devices support mixed-precision tensor core operations. Recent works are actively exploring activation quantization (lutkevich2024_quarot; liu2025_spinquant; liu2024llmqat), either through co-design during training or post-training calibration, and integrating such methods with BLR could further reduce overheads. Finally, while our experiments on billion-parameter models were constrained to partial re-training (only 400400-$4000$ training steps) due to limited compute resources, extended re-training, as was feasible for smaller models ($10^{6}$ training steps), could further narrow the accuracy gap to dense baselines.
@@ A.3 Experimental Details
We employed the Llama-3.2 (llama3) model with 1.24B parameters as one of the representative medium-sized models. The model was compressed by 50% using low-rank, Monarch, and BLAST weight parameterizations. Specifically, we replaced the query and output projection weights in the attention modules, as well as the up projection, gate projection, and down projection weights in the feed-forward network modules.
The rank and number of blocks used in each experiment are reported in Table 3. For low-rank compression, the weights were factorized via singular value decomposition (SVD). For Monarch compression, we applied block-wise SVD, where each block had rank $r^{\prime}=\tfrac{r}{b}$. Following (blast-neurips24), BLAST compression was obtained by applying 300 steps of preconditioned gradient descent to factorize the weights into BLAST factors. All three methods reduced the model size to 0.6B parameters.
As in (blast-neurips24), the compressed models were re-trained. Specifically, the compressed weights were fine-tuned for 4000 steps on a subset [^2] https://huggingface.co/datasets/DKYoon/SlimPajama-6B of the SlimPajama dataset (slimpajama-dataset), using a learning rate of $8\times 10^{-4}$ with linear decay scheduling.

### v2

@@ 6 Conclusion
Our optimized BLAST and Monarch routines still lag behind low-rank decompositions in terms of speed, a limitation rooted in their blocked structure, which generates more intermediate outputs. This overhead, however, could be mitigated by future techniques such as intermediate activation quantization, provided accuracy is preserved and target devices support mixed-precision tensor core operations. Recent works are actively exploring activation quantization (Ashkboos et al., 2024; Liu et al., 2025, 2024), either through co-design during training or post-training calibration, and integrating such methods with BLR could further reduce overheads. Finally, while our experiments on billion-parameter models were constrained to partial re-training (only 400400-$4000$ training steps) due to limited compute resources, extended re-training, as was feasible for smaller models ($10^{6}$ training steps), could further narrow the accuracy gap to dense baselines.
@@ A.3 Experimental Details
We employed the Llama-3.2 (Grattafiori et al., 2024) model with 1.24B parameters as one of the representative medium-sized models. The model was compressed by 50% using low-rank, Monarch, and BLAST weight parameterizations. Specifically, we replaced the query and output projection weights in the attention modules, as well as the up projection, gate projection, and down projection weights in the feed-forward network modules.
The rank and number of blocks used in each experiment are reported in Table 3. For low-rank compression, the weights were factorized via singular value decomposition (SVD). For Monarch compression, we applied block-wise SVD, where each block had rank $r^{\prime}=\tfrac{r}{b}$. Following (Lee et al., 2024), BLAST compression was obtained by applying 300 steps of preconditioned gradient descent to factorize the weights into BLAST factors. All three methods reduced the model size to 0.6B parameters.
As in (Lee et al., 2024), the compressed models were re-trained. Specifically, the compressed weights were fine-tuned for 4000 steps on a subset [^2] https://huggingface.co/datasets/DKYoon/SlimPajama-6B of the SlimPajama dataset (Soboleva et al., 2023), using a learning rate of $8\times 10^{-4}$ with linear decay scheduling.